What number is described by the facts below?
Fact: The number is an integer.
Fact: The cube of the number is less than 50,000.
Fact: The square of the number is greater than 1,250.

Answer by rwm(914)

36
x^3< 50000
x^2>1250
35^2=1225
36^3=46656
37^3> 50000
37^3=50653
so an integer between 37 and 35